Hello, Can someone show me how to work this problem. I'm trying and cannot seem to figure this out. Thanks in Advance.
The Web-based company Oh Baby! Gifts has a goal of processing 95 percent of its orders on the same day they are received. If 485 out of the next 500 orders are processed on the same day, would this prove that they are exceeding their goal, using a = .025? (See story.news.yahoo.com accessed June 25, 2004.)

Answer by stanbon(75887)

The Web-based company Oh Baby! Gifts has a goal of processing 95 percent of its orders on the same day they are received. If 485 out of the next 500 orders are processed on the same day, would this prove that they are exceeding their goal, using a = .025?
---------------------------
Ho: p = 0.95
Ha: p > 0.95
---------------
p-hat = 485/500 = 0.97
--------------
Critical value for one-tail z test with alpha = 0.025 = 1.96
--------------
Test statistic:
z(0.97) = (0.97-0.95)*sqrt[0.95*0.05/500] = 2.052
---------------
Since the test statistic is greater than the critical value,
reject Ho.
Conclusion: The test provides statistical evidence they are exceeding their goal.
====================
Cheers,
Stan H.
====================